OpenStudy (anonymous):

solve the system by the elimination method. 5x+3y=-11 7x-2y=17

OpenStudy (anonymous):

(5x+3y=-11)*2 ------- 10x+6y=-22 (7x-2y=17)*3 ----- 21x-6y=51 add the equations 10x+6y+21x-6y=-22+51 31x=29 x=29/31 i think....

OpenStudy (anonymous):

is this as an ordered pair. simplify the answer.